The value of x-yx-y when x=2 & y= -2
 

Grade:10

2 Answers

Ravi
9 years ago
Substitute the values of x and y in the expression
2-(-2)2-(-2)
=2-(-2)2+2
=2-(-2)4
=2-(16)
=2-16
=-14
.....................................................................................................................................
parthasarathy
9 years ago
-14